OpenStudy (anonymous):

Simplify using the distributive property. 5x + 3(x + 7) = 11 - (2x + 5) How do you figure out 11-(2x+5)? That is what I am having trouble remembering..

OpenStudy (anonymous):

@iGreen @iambatman @CausticSyndicalist @xo.minnie.xox @AdmiralAngel25

OpenStudy (studygurl14):

5x + 3(x + 7) = 11 - (2x + 5) 5x +3x + 21 = 11 -2x - 5 8x + 21 = -2x + 6 Can you do the rest?

OpenStudy (anonymous):

@StudyGurl14 oh, ok, I understand the first part, it is the 11-(2x+5) part that stumps me, so for that, let's say it is 12-(3x+7), you would just make the 3x a -3x right?

OpenStudy (studygurl14):

in that instance, the 3x would become -3x, adn the 7 would become -7

OpenStudy (anonymous):

@StudyGurl14 ok, thanks so much! :) Do you want a fan? I will give you a medal.

OpenStudy (studygurl14):

Could I have a fan + tesitmonial please? It would be greatly appreciated!

OpenStudy (studygurl14):

And I'll give you a medal too

OpenStudy (anonymous):

@StudyGurl14 haha no problem! You're sweet ;) And smart ;)